OpenStudy (anonymous):

Simplify the expression. 6(6) + 6(4) A. 60 B. 10 C. 16 D. 12

OpenStudy (anonymous):

36+24 = 60 You really need to work on figuring these kinds of problems out. This is just basic math.

OpenStudy (adi3):

6(6) = 36 6(4) = 24 then u add them together 36 + 24 = 60.

OpenStudy (ilovewolf):

6(6)=6*6=36 6(4)=6*4=24 36+24=60 c:

OpenStudy (anonymous):

@MathNinja_C thank you for the help. and I know its basic math clearly if I didn't need help I wouldn't have asked I'm not very good with math. JA

OpenStudy (anonymous):

If you can communicate by typing complete sentences into a computer then you have the ability to add and subtract numbers.

OpenStudy (ilovewolf):

Harsh......